OpenStudy (ciel21):

If the quotient is 1 and the dividend is 4, what is the divisor?

OpenStudy (bibby):

the dividend is the thing outside of the division.

OpenStudy (ciel21):

thanks ^w^

OpenStudy (bibby):

o3o

OpenStudy (ciel21):

i may have a couple more idk >n<

OpenStudy (bibby):

no worries. I'm eating dinner so I have all the time in the world

OpenStudy (ciel21):

okay ^w^

OpenStudy (ciel21):

in the school cafeteria, each table can have 4 chairs. If there are 8 chairs, how many tables will have chairs?

OpenStudy (ciel21):

wrong question

OpenStudy (ciel21):

>~<

OpenStudy (bibby):

I'll just wait for the right one then lol

OpenStudy (ciel21):

Choose the problem that correctly answers this question. At summer camp, there are four canoes. Twelve students have enrolled in the canoeing class. How many life preservers will be placed in each canoe if each canoe carries the same number of students? (Points : 4) a. 12/3=12 b.4+12=16 c.12/6=2 d.12/4=3

OpenStudy (bibby):

You know that you're dividing because you have to give out preservers to each boat. 12 preservers, 4 boats. That means each one gets 3. the "carries the same number of students" is another hint at division

OpenStudy (ciel21):

oh >w< thanks

OpenStudy (ciel21):

thats it thanks ^w^

OpenStudy (bibby):

yay, no problem :D